Your Uncle gives you $7,000 today. You want to put it away and use it as a down payment on a new car in 2 years. You put it in a mutual fund investment with an online discount broker, earning 5% per year.
a) (6 points) What will you have at the end of 2 years, assuming annual compounding? SHOW ALL WORK FOR FULL CREDIT using the TI BAII Plus Calculator.
b) (2 points) What is the total interest in $ earned on this investment over the 2 years? Show work.
c) (4 points) If the investment account compounds interest monthly, what will you have at the end of 2 years? SHOW ALL WORK FOR FULL CREDIT using the TI BAII Plus Calculator.

Please ensure to show each button that needs to be pressed on the TI BA II Plus Calculator.
Answer #1

1.
N=2
PMT=0
PV=-7000
I/Y=5%
CPT FV=7717.50

Amount accumulated=7717.50

2.
=7717.50-7000=717.50

3.
N=2*12=24
PMT=0
PV=-7000
I/Y=5%/12
CPT FV=7734.59

Amount accumulated=7734.59